The Gauteng traffic department records show that 25% of all drivers wear seatbelts. In a random sample of 400 cars stopped at a roadblock in Gauteng, 152 of the drivers were wearing seatbelts. A 90% confidence interval for the proportion in the population who wear seatbelts is given by (round final answer to two decimal places):



1
Expert's answer
2021-10-12T09:16:48-0400

"p=0.25 \\

n=400 \\

\hat{p} = \frac{152}{400}=0.38 \\

CI= \hat{p}±Z^* \times \sqrt{ \frac{\hat{p}(1- \hat{p}}{n} }"

"Z^*= 1.645" (for 90 % confidence interval)

"CI = 0.38 ± 1.645 \times \sqrt{\frac{0.38 \times 0.62}{400}} \\

= 0.38± 0.04 \\

= (0.34, 0.42)"